OpenStudy (roadjester):

int_{?}^{?}int_{?}^{?}_{D} xdA , D = forall ({x,y})in 0le x le pi, 0le y le sinx

OpenStudy (anonymous):

i think some syntax needs fixing. :D

OpenStudy (roadjester):

\[\int\limits_{?}^{?}\int\limits_{?}^{?}_{D} xdA , D = \forall ({x,y})\in 0\le x \le \pi, 0\le y \le sinx\]

OpenStudy (turingtest):

\[\int\int\limits_D xdA\]\[\{D:\forall(x,y)\in0\le x\le\pi,0\le y\le\sin x\}\]

OpenStudy (roadjester):

ok, thanks Turning test

OpenStudy (roadjester):

question, how did you get rid of the "?"

OpenStudy (anonymous):

dA = area, so in this case it is dxdy or dydx (depends on how you want to integrate). The inequalities are your bounds of integration. just choose your bounds on the inner most integral for however you setup your dxdy or dydx. DOes that make sense?

OpenStudy (turingtest):

I just typed the latex directly, but if you use the equation editor you can just delete the {?} stuff I just wrote \int for\[\int\]

OpenStudy (experimentx):